An old go-kart with a mass of 300 kg is traveling in a straight line at 80 m/s. It is
followed by a 4-wheeler with mass of 200 kg moving at 60 m/s. How fast is the center
of mass moving?
1
Expert's answer
2020-12-06T17:19:50-0500
"v_{cm}=\frac{m_1v_1+m_2v_2}{m_1+m_2}\\
v_{cm}=\frac{300(80)+200(60)}{300+200}=72\frac{m}{s}"